OpenStudy (anonymous):

find f(n(x)) when f(x)=x-3 and n(x)=3x-5

hartnn (hartnn):

to get f(n(x)), just replace x in f(x) by n(x)=3x-5

OpenStudy (anonymous):

So, 3(x-3)-5?

hartnn (hartnn):

nopes, what u did is reverse, u put x in n(x) by f(x)=x-3 you had to put 'x' in f(x) by 3x-5

OpenStudy (anonymous):

So, (3x-5)-3

hartnn (hartnn):

yes, that is correct f(n(x))

hartnn (hartnn):

and it can be simplified

OpenStudy (anonymous):

3x-8

hartnn (hartnn):

thats correct final answer :)

OpenStudy (anonymous):

can you help me with one more?

hartnn (hartnn):

sure

OpenStudy (anonymous):

k(h(x)) when k(x)=x\[^{2}\]+4 and h(x)=x+2

OpenStudy (anonymous):

when K(x)=x^2 +4 and h(x)=x+2

hartnn (hartnn):

to get k(h(x)), just replace 'x' in k(x) by h(x) = x+2

hartnn (hartnn):

wherever u see x in k(x) , replace it by x+2

OpenStudy (anonymous):

x+2+2, x+4?

hartnn (hartnn):

nopes,